# Low-Altitude ISAC with Rotatable Active and Passive Arrays

### v1

@@ III-C3 Overall algorithm and complexity analysis for subproblem 2
III-D Subproblem 3: Array Rotation Angles
Given $\boldsymbol{W}$ and $\boldsymbol{\theta}$ from the previous subproblems, we jointly optimize the array rotations $(\boldsymbol{r}^{\mathrm{B}},\boldsymbol{r}^{\mathrm{R}})$ of the BS and RIS via
$\text{(P4)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\right)}$
$-\rho\max\{\text{MSE}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)-\eta_{\text{th}},0\}\!\!\!$ (67)
s.t. $\left(22\text{e}\right),\left(22\text{f}\right)$
@@ III-D1 Gradient computation
The key step of PGD is to derive the gradient expression for the complex objective function (67). In the following, we consider the case that the rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})^{(\ell)}$ in the current iteration violate the MSE threshold constraint. Thus, the MSE term is involved in the cost function; otherwise, the cost and gradient functions relate only to the communication sum rate. We adopt $\boldsymbol{r}^{\text{A}}$ for any array rotation variable with $\text{A}\in\{\text{B,R}\}$ for brevity as defined in (2). Further, define a six-dimensional vector $\boldsymbol{r}$ for all rotations as below
$\boldsymbol{r}\triangleq[(\boldsymbol{r}^{\text{B}})^{T},(\boldsymbol{r}^{\text{R}})^{T}]^{T}=\left[r_{x}^{\text{B}},r_{y}^{\text{B}},r_{z}^{\text{B}},r_{x}^{\text{R}},r_{y}^{\text{R}},r_{z}^{\text{R}}\right]^{T}\in\mathbb{R}^{6}.$ (68)
The gradient of (67) w.r.t. the generic rotation vector $\boldsymbol{r}$ is
$\frac{\partial F\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}=$ $-\frac{2\rho}{A}\sum_{a=1}^{A}{\left(\mathcal{P}_{a}\left(\boldsymbol{r}\right)-\mathcal{P}_{\text{d},a}\right)\frac{\partial\mathcal{P}_{a}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}}$
@@
where $\frac{\partial\text{SINR}_{k}(\boldsymbol{r})}{\partial\boldsymbol{r}}$ and $\frac{\partial\mathcal{P}_{a}(\boldsymbol{r})}{\partial\boldsymbol{r}}$ are the two main computational branches.
Define the signal and interference terms in the SINR expression by $C_{1,k}\left(\boldsymbol{r}\right)=|\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}\right)\boldsymbol{w}_{k}|^{2}$ and $C_{2,k}\left(\boldsymbol{r}\right)=\sum_{i\neq k}{|}\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}\right)\boldsymbol{w}_{i}|^{2}+\sigma_{k}^{2}$, respectively. Then we have
@@
$\frac{\partial\boldsymbol{d}_{v}^{\text{A}}(\boldsymbol{r}^{\text{A}})}{\partial\boldsymbol{r}^{\text{A}}}\!=\!\Big[\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{x}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{y}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{z}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}}\Big].\!\!\!$ (74c)
The derivatives of the rotation matrix $\boldsymbol{R}\left(\boldsymbol{r}^{\text{A}}\right)=\boldsymbol{R}_{x}\left(r_{x}^{\text{A}}\right)\boldsymbol{R}_{y}\left(r_{y}^{\text{A}}\right)\boldsymbol{R}_{z}\left(r_{z}^{\text{A}}\right)$ w.r.t each rotation variable are
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{x}^{\text{A}}}=\frac{\partial\boldsymbol{R}_{x}(r_{x}^{\text{A}})}{\partial r_{x}^{\text{A}}}\boldsymbol{R}_{y}(r_{y}^{\text{A}})\boldsymbol{R}_{z}(r_{z}^{\text{A}}),$ (75a)
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{y}^{\text{A}}}=\boldsymbol{R}_{x}(r_{x}^{\text{A}})\frac{\partial\boldsymbol{R}_{y}(r_{y}^{\text{A}})}{\partial r_{y}^{\text{A}}}\boldsymbol{R}_{z}(r_{z}^{\text{A}}),$ (75b)
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{z}^{\text{A}}}=\boldsymbol{R}_{x}(r_{x}^{\text{A}})\boldsymbol{R}_{y}(r_{y}^{\text{A}})\frac{\partial\boldsymbol{R}_{z}(r_{z}^{\text{A}})}{\partial r_{z}^{\text{A}}},$ (75c)
@@
$\frac{\partial\boldsymbol{R}_{x}\left(r_{x}^{\text{A}}\right)}{\partial r_{x}^{\text{A}}}=\left[\begin{matrix}0&0&0\\ 0&-\sin r_{x}^{\text{A}}&-\cos r_{x}^{\text{A}}\\ 0&\cos r_{x}^{\text{A}}&-\sin r_{x}^{\text{A}}\\ \end{matrix}\right],$ (76a)
$\frac{\partial\boldsymbol{R}_{y}\left(r_{y}^{\text{A}}\right)}{\partial r_{y}^{\text{A}}}=\left[\begin{matrix}-\sin r_{y}^{\text{A}}&0&\cos r_{y}^{\text{A}}\\ 0&0&0\\ -\cos r_{y}^{\text{A}}&0&-\sin r_{y}^{\text{A}}\\ \end{matrix}\right],$ (76b)
$\frac{\partial\boldsymbol{R}_{z}\left(r_{z}^{\text{A}}\right)}{\partial r_{z}^{\text{A}}}=\left[\begin{matrix}-\sin r_{z}^{\text{A}}&-\cos r_{z}^{\text{A}}&0\\ \cos r_{z}^{\text{A}}&-\sin r_{z}^{\text{A}}&0\\ 0&0&0\\ \end{matrix}\right].$ (76c)
@@
$\tfrac{\partial\sqrt{G}}{\partial\boldsymbol{r}}=\frac{1}{2\sqrt{G}}\frac{\partial G}{\partial\boldsymbol{r}}.$ (77a)
$\frac{\partial G^{\text{A}}}{\partial\boldsymbol{r}^{\text{A}}}=G_{0}^{\text{A}}\,p^{\text{A}}\big(-\boldsymbol{n}^{T}\boldsymbol{u}\big)^{p^{\text{A}}-1}\Big(-\frac{\partial\boldsymbol{n}^{T}}{\partial\boldsymbol{r}^{\text{A}}}\boldsymbol{u}\Big),$ (77b)
@@
$\,\,\frac{\partial\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{l=1}^{L_{k}}{a_{k,l}\bigg[\frac{\partial\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})+\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}\bigg]}$ (78a)
$\,\,\frac{\partial\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{q=1}^{Q_{k}}{c_{k,q}\bigg[\frac{\partial\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{k,l}^{\text{RU}}(\boldsymbol{r}^{\text{R}})+\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\boldsymbol{t}_{k,l}^{\text{RU}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\bigg]}.$ (78b)
$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{p=1}^{P}{b_{p}}\bigg[\frac{\partial\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\sqrt{G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\partial\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}\bigg],$ (78c)
$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{p=1}^{P}{b_{p}}\bigg[\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\sqrt{G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\partial\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}}(\boldsymbol{r}^{\text{B}}))(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{k,p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{k,p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})\Big(\frac{\partial\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\Big)^{H}\bigg],$ (78d)
@@
According to the sensing beam pattern (18) and the sensing channel model (14), for the beam-pattern derivative,
$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\text{A}}}=2\Re\Big\{\Big(\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\mathrm{A}}}\Big)^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})\Big\},$ (79)
where $\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\mathrm{A}}}$ is computed in (80).
$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}=a_{a}\frac{\partial\left(\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}\boldsymbol{t}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)\right)}{\partial\boldsymbol{r}^{\text{B}}}+c_{a}\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right),$ (80a)
$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}=a_{a}\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}}\boldsymbol{t}^{\text{RT}}+c_{a}\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\mathrm{diag}\left(\boldsymbol{\theta}\right)\frac{\partial\left(\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right)\right)}{\partial\boldsymbol{r}^{\text{R}}}$ (80b)
@@
It is also convenient to gather partials into Jacobians $\boldsymbol{J}_{\text{S},a}^{(\text{A})}\triangleq\left[\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{x}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{y}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{z}^{\text{A}}}\right]\in\mathbb{C}^{N_{t}\times 3}$ similar to (74), so that (79) compacts to
$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\mathrm{A}}}=2\,\Re\!\Big\{\boldsymbol{J}_{\text{S},a}^{(\text{A})\,H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\Big\}\in\mathbb{R}^{3}.$ (81)

### v2

@@ III-C3 Overall algorithm and complexity analysis for subproblem 2
III-D Subproblem 3: Array Rotations
Given $\boldsymbol{W}$ and $\boldsymbol{\theta}$ from the previous subproblems, we jointly optimize the array rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ of the BS and RIS via
$\text{(P4)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\right)}-\rho\text{NMSE}(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})\!\!\!$ (68)
[…]
@@ III-D1 Gradient computation
The key step of PGD is to derive the gradient expression for the complex objective function (68). We adopt $\boldsymbol{r}^{\text{A}}$ for any array rotation variable with $\text{A}\in\{\text{B,R}\}$ for brevity as defined in (2). Further, define a stacked six-dimensional vector $\boldsymbol{r}$ for all rotations as below
$\boldsymbol{r}\triangleq[(\boldsymbol{r}^{\text{B}})^{T},(\boldsymbol{r}^{\text{R}})^{T}]^{T}=\left[r_{\mathrm{x}}^{\text{B}},r_{\mathrm{y}}^{\text{B}},r_{\mathrm{z}}^{\text{B}},r_{\mathrm{x}}^{\text{R}},r_{\mathrm{y}}^{\text{R}},r_{\mathrm{z}}^{\text{R}}\right]^{T}\in\mathbb{R}^{6}.$ (69)
The gradient of (68) w.r.t. the stacked rotation vector $\boldsymbol{r}$ is
$\frac{\partial F\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}=$ $-\frac{2\rho}{D}\sum_{a=1}^{A}{\left(\mathcal{P}_{a}\left(\boldsymbol{r}\right)-\bar{\mathcal{P}}_{\text{d},a}\right)\frac{\partial\mathcal{P}_{a}\left(\boldsymbol{r}\right)}{\partial\boldsymbol{r}}}$
@@
where $\frac{\partial\text{SINR}_{k}(\boldsymbol{r})}{\partial\boldsymbol{r}}$ and $\frac{\partial\mathcal{P}_{a}(\boldsymbol{r})}{\partial\boldsymbol{r}}$ are the two main computational branches. Define the signal and interference terms in the SINR expression by $C_{1,k}\left(\boldsymbol{r}\right)=|\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}\right)\boldsymbol{w}_{k}|^{2}$ and $C_{2,k}\left(\boldsymbol{r}\right)=\sum_{i\neq k}{|}\boldsymbol{f}_{k}^{H}\left(\boldsymbol{r}\right)\boldsymbol{w}_{i}|^{2}+\sigma_{k}^{2}$, respectively. Then we have
@@
$\frac{\partial\boldsymbol{d}_{v}^{\text{A}}(\boldsymbol{r}^{\text{A}})}{\partial\boldsymbol{r}^{\text{A}}}\!=\!\Big[\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{x}}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{y}}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}},\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{z}}^{\text{A}}}\overline{\boldsymbol{d}}_{v}^{\text{A}}\Big].\!\!\!$ (75c)
The derivatives of the rotation matrix $\boldsymbol{R}\left(\boldsymbol{r}^{\text{A}}\right)=\boldsymbol{R}_{\mathrm{x}}\left(r_{\mathrm{x}}^{\text{A}}\right)\boldsymbol{R}_{\mathrm{y}}\left(r_{\mathrm{y}}^{\text{A}}\right)\boldsymbol{R}_{\mathrm{z}}\left(r_{\mathrm{z}}^{\text{A}}\right)$ w.r.t each rotation variable are
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{x}}^{\text{A}}}=\frac{\partial\boldsymbol{R}_{\mathrm{x}}(r_{\mathrm{x}}^{\text{A}})}{\partial r_{\mathrm{x}}^{\text{A}}}\boldsymbol{R}_{\mathrm{y}}(r_{\mathrm{y}}^{\text{A}})\boldsymbol{R}_{\mathrm{z}}(r_{\mathrm{z}}^{\text{A}}),$ (76a)
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{y}}^{\text{A}}}=\boldsymbol{R}_{\mathrm{x}}(r_{\mathrm{x}}^{\text{A}})\frac{\partial\boldsymbol{R}_{\mathrm{y}}(r_{\mathrm{y}}^{\text{A}})}{\partial r_{\mathrm{y}}^{\text{A}}}\boldsymbol{R}_{\mathrm{z}}(r_{\mathrm{z}}^{\text{A}}),$ (76b)
$\frac{\partial\boldsymbol{R}(\boldsymbol{r}^{\text{A}})}{\partial r_{\mathrm{z}}^{\text{A}}}=\boldsymbol{R}_{\mathrm{x}}(r_{\mathrm{x}}^{\text{A}})\boldsymbol{R}_{\mathrm{y}}(r_{\mathrm{y}}^{\text{A}})\frac{\partial\boldsymbol{R}_{\mathrm{z}}(r_{\mathrm{z}}^{\text{A}})}{\partial r_{\mathrm{z}}^{\text{A}}},$ (76c)
@@
$\frac{\partial\boldsymbol{R}_{\mathrm{x}}\left(r_{\mathrm{x}}^{\text{A}}\right)}{\partial r_{\mathrm{x}}^{\text{A}}}=\left[\begin{matrix}0&0&0\\ 0&-\sin r_{\mathrm{x}}^{\text{A}}&-\cos r_{\mathrm{x}}^{\text{A}}\\ 0&\cos r_{\mathrm{x}}^{\text{A}}&-\sin r_{\mathrm{x}}^{\text{A}}\\ \end{matrix}\right],$ (77a)
$\frac{\partial\boldsymbol{R}_{\mathrm{y}}\left(r_{\mathrm{y}}^{\text{A}}\right)}{\partial r_{\mathrm{y}}^{\text{A}}}=\left[\begin{matrix}-\sin r_{\mathrm{y}}^{\text{A}}&0&\cos r_{\mathrm{y}}^{\text{A}}\\ 0&0&0\\ -\cos r_{\mathrm{y}}^{\text{A}}&0&-\sin r_{\mathrm{y}}^{\text{A}}\\ \end{matrix}\right],$ (77b)
$\frac{\partial\boldsymbol{R}_{\mathrm{z}}\left(r_{\mathrm{z}}^{\text{A}}\right)}{\partial r_{\mathrm{z}}^{\text{A}}}=\left[\begin{matrix}-\sin r_{\mathrm{z}}^{\text{A}}&-\cos r_{\mathrm{z}}^{\text{A}}&0\\ \cos r_{\mathrm{z}}^{\text{A}}&-\sin r_{\mathrm{z}}^{\text{A}}&0\\ 0&0&0\\ \end{matrix}\right].$ (77c)
@@
$\frac{\partial\sqrt{G^{\text{A}}}}{\partial\boldsymbol{r}^{\text{A}}}=\frac{1}{2\sqrt{G^{\text{A}}}}\frac{\partial G^{\text{A}}}{\partial\boldsymbol{r}^{\text{A}}},$ (78a)
[…]
@@
$\,\,\frac{\partial\boldsymbol{h}_{k}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{l=1}^{L_{k}}{\delta_{k,l}\bigg[\frac{\partial\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})+\sqrt{G_{k,l}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\boldsymbol{t}_{k,l}^{\text{BU}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}\bigg]},$ (79a)
$\,\,\frac{\partial\boldsymbol{g}_{k}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{q=1}^{Q_{k}}{\psi_{k,q}\bigg[\frac{\partial\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\partial\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{R}})+\sqrt{G_{k,q}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\boldsymbol{t}_{k,q}^{\text{RU}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\bigg]},$ (79b)
$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{B}}}=\sum_{p=1}^{P}{\gamma_{p}}\bigg[\frac{\partial\sqrt{G_{p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}}{\partial\boldsymbol{r}^{\text{B}}}\sqrt{G_{p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\frac{\partial\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})}{\partial\boldsymbol{r}^{\text{B}}}(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}\bigg],$ (79c)
$\,\,\frac{\partial\boldsymbol{B}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{R}}}=\sum_{p=1}^{P}{\gamma_{p}}\bigg[\sqrt{G_{p}^{\text{B}}(\boldsymbol{r}^{\text{B}})}\frac{\partial\sqrt{G_{p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}}{\partial\boldsymbol{r}^{\text{R}}}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}}(\boldsymbol{r}^{\text{B}}))(\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}}))^{H}+\sqrt{G_{p}^{\text{B}}(\boldsymbol{r}^{\text{B}})G_{p}^{\text{R}}(\boldsymbol{r}^{\text{R}})}\boldsymbol{t}_{p}^{\text{BR}}(\boldsymbol{r}^{\text{B}})\Big(\frac{\partial\boldsymbol{t}_{p}^{\text{RB}}(\boldsymbol{r}^{\text{R}})}{\partial\boldsymbol{r}^{\text{R}}}\Big)^{H}\bigg],$ (79d)
@@
According to the sensing beampattern (18) and the sensing channel model (II-C), for the beampattern derivative,
$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\text{A}}}=2\Re\Big\{\Big(\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{A}}}\Big)^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})\Big\},$ (80)
where $\frac{\partial\boldsymbol{f}_{\text{S},a}(\boldsymbol{r})}{\partial\boldsymbol{r}^{\text{A}}}$ is computed in (81).
$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}=\frac{\partial\left(\sqrt{G_{a}^{\text{B}}\left(\boldsymbol{r}^{\text{B}}\right)}\boldsymbol{t}^{\text{BT}}\left(\boldsymbol{r}^{\text{B}}\right)\right)}{\partial\boldsymbol{r}^{\text{B}}}+\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{B}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right),$ (81a)
$\frac{\partial\boldsymbol{f}_{\text{S},a}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}=\frac{\partial\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)}{\partial\boldsymbol{r}^{\text{R}}}\mathrm{diag}\left(\boldsymbol{\theta}\right)\sqrt{G_{a}^{\text{R}}}\boldsymbol{t}^{\text{RT}}+\boldsymbol{B}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\right)\mathrm{diag}\left(\boldsymbol{\theta}\right)\frac{\partial\left(\sqrt{G_{a}^{\text{R}}\left(\boldsymbol{r}^{\text{R}}\right)}\,\boldsymbol{t}^{\text{RT}}\left(\boldsymbol{r}^{\text{R}}\right)\right)}{\partial\boldsymbol{r}^{\text{R}}}$ (81b)
@@
It is also convenient to gather partials into Jacobians $\boldsymbol{J}_{\text{S},a}^{(\text{A})}\triangleq\left[\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{\mathrm{x}}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{\mathrm{y}}^{\text{A}}},\frac{\partial\boldsymbol{f}_{\text{S},a}}{\partial r_{\mathrm{z}}^{\text{A}}}\right]\in\mathbb{C}^{M\times 3}$ similar to (75), so that (80) compacts to
$\frac{\partial\mathcal{P}_{a}}{\partial\boldsymbol{r}^{\text{A}}}=2\,\Re\!\Big\{\boldsymbol{J}_{\text{S},a}^{(\text{A})\,H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\Big\}\in\mathbb{R}^{3}.$ (82)